OpenStudy (sleepyjess):

Confused on how to do this: Find the angle between u = <-3, 2> and v = <2, 3>

OpenStudy (sleepyjess):

@Loser66

OpenStudy (loser66):

you have formula for this, it is cos (u and v) =.....

OpenStudy (sleepyjess):

I know \(\sf cos\theta = \dfrac{u*v}{||u||~||v||}\), but I'm not really sure how to use that.

OpenStudy (loser66):

so, u dot v, not u*v, what do you get for u dot v(the numerator)

OpenStudy (sleepyjess):

-3 * 2 + 2*3 -6 + 6 -

OpenStudy (sleepyjess):

*0

OpenStudy (anonymous):

Well, take arccos of both sides. \[ \theta = \arccos\left(\frac{\mathbf u\cdot \mathbf v}{\|\mathbf u\|\|\mathbf v\|}\right) \]

OpenStudy (loser66):

so, the whole thing is 0, cos (what?) = 0?

OpenStudy (sleepyjess):

cos(90) = 0

OpenStudy (sleepyjess):

but what about the denominator of the fraction? That is what is confusing me

OpenStudy (loser66):

DAT SIT.

OpenStudy (anonymous):

\[ \|\mathbf u\| = \sqrt{\mathbf u \cdot \mathbf u} \]

OpenStudy (kainui):

Maybe this helps, just imagine the vector as a hypotenuse of a right triangle, so we can use the pythagorean theorem to get its length. \[\Large \Large v = \langle x,y,z \rangle \\ \Large ||v|| = \sqrt{x^2+y^2+z^2}\]

OpenStudy (sleepyjess):

Ok, so the denominator would be \(\sqrt{-3^2 + 2^2}\sqrt{2^2 + 3^2}\)?

OpenStudy (anonymous):

\[||u|| = \sqrt{(-3)^2+(2)^2}\]

OpenStudy (anonymous):

But yes, you got the right idea

OpenStudy (sleepyjess):

ok ^_^ thank you all

OpenStudy (anonymous):

That's one messed up triangle

OpenStudy (sleepyjess):

How does this relate to a triangle? *confused*

OpenStudy (kainui):

All vectors are just the hypotenuse of triangles.